Prev Next

SW&IT Meeting Minutes 2018-03-26

Attendees: David, Doug, Igor, Joe, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Previous Action items

  • TCS/OT binocular shared code, view in github (JA) ONGOING
  • SWG Summer Shutdown 2018 plan - follow up 'need more info' items, OC / Laura engineering plan (MB) ONGOING
  • swg management - explore new tool https://github.com/LBTO/issues/projects (All)
  • github watch (All)
  • svn survey follow up (MB)
  • wiki report guideline, privacy?, follow up (MB) DONE
  • follow up NetBrain (MB) DONE
  • follow up VM SSD (MB) DONE

Outlook

Issues


Issue Number Status Priority Subject Assigned To Next Action Entered By Closed By Last Activity By Next Actioned By Submitted By Reopened By
3949 Open Medium Rotator time-to-limit predictions are inaccurate swadmin pkubanek sallanson   pkubanek Admin, Software sallanson  
6138 Open Medium LUCI stops executing a script swadmin apramskiy tsargent   isola Admin, Instrument tsargent  
6565 Open Medium AO: In UAO, Adsec arbitrator goes offline if a command is sent during high safe-skip % iqadmin xzhang avaz   spalding Admin, IQ avaz  
6758 Open Medium 4D upgrade/New PC iqadmin gbrusa xzhang   mbec Admin, IQ xzhang  
7099 Open High Core-A damage, water incursion in CRA swadmin shooper shooper   shooper Hooper, Stephen shooper  
7134 Closed Medium Synology Rackstation1 failure, March 2018 swadmin edmeades edmeades edmeades edmeades Bec, Matthieu edmeades  
7141 Open Low AOS stopAO command triggers Ice exceptions, transient alarms swadmin xzhang sallanson   xzhang Bec, Matthieu sallanson  
7143 Open Medium MODS ROI command for imaging swadmin okuhn Jpower   okuhn Bec, Matthieu Jpower  
7146 Closed High Mountain Synology Passive server problems swadmin shooper shooper shooper edmeades Hooper, Stephen shooper  
7150 Open Medium Local VM storage install swadmin edmeades edmeades   edmeades Bec, Matthieu edmeades  
7170 Open Medium OVMS+ sensitivities calculated in arcsec but should be mas   ksummers ksummers   ksummers Summers, Kellee ksummers  
7171 Open High oac Tucson machine is unreachable swadmin shooper ksummers   mbec Bec, Matthieu ksummers  


Issue Number Status PrioritySorted ascending Subject Assigned To Next Action Entered By Closed By Last Activity By Next Actioned By Submitted By Reopened By
7023 Open Critical MODS2 Red Head Electronics Box problem instadmin rpogge okuhn   okuhn Reynolds, Robert okuhn  
6255 Open High AO: DX AdSec logs error: FAILED PACKET FROM BCU 2 - CMD TIMEOUT CA iqadmin gbrusa dlmiller   spalding Miller, Doug dlmiller  
7138 Closed High Utility Chiller CH0002 circuit one failed woadmin woadmin phartley phartley phartley Hartley, Patrick phartley  
7168 Open High MODS2 Blue Race Condition, IMCS Lock Starts During Readout of Previous Exposure instadmin rpogge brothberg   okuhn Kuhn, Olga brothberg  
7164 Open Low TPLC Heartbeat Bad engadmin   gbechetti   gbechetti   gbechetti  
7165 Closed Low 480v heater on 3L     cbolyard cbolyard cbolyard   cbolyard  
6304 Open Medium AdSec arbitrator goes offline iqadmin puglisi sallanson   spalding Admin, IQ sallanson  
7009 Closed Medium no flux from MODS2 VFLAT lamp instadmin esolheid Jpower okuhn okuhn Reynolds, Robert Jpower  
7130 Open Medium SX and DX RIPs associated with BCUCURRENTVCCP errors iqadmin   spalding   spalding   spalding  
7136 Closed Medium DXM1 Actuator 233 engadmin   sallanson mgardiner mgardiner   sallanson  
7145 Open Medium some MODS headers have duplicate keywords instadmin jmason okuhn   okuhn Reynolds, Robert okuhn  
7159 Closed Medium stripes along dispersion axis instadmin rpogge okuhn rpogge okuhn Reynolds, Robert okuhn  
7161 Open Medium 4D bracket beside DX PFU blocking access engadmin woadmin ajarvinen   engadmin Admin, Engineering ajarvinen  
7162 Open Medium ARGOS: Arbitrator GUI software updates. engadmin   acardwell   acardwell   acardwell  
7163 Open Medium 4D installation for SX engadmin jkern jkern   engadmin Admin, Engineering jkern  
7166 Closed Medium DX M1 voltage failure and air pressure loop out of range engadmin   dhuerta jhill jhill   dhuerta  
7167 Open Medium PEPSI solid models instadmin engadmin jkern   jkern Kern, Jonathan jkern  
7169 Open Monitor MODS-1 Blue IMCS Lock Fails Repeatedly instadmin   brothberg   okuhn   brothberg  

Team reports

AC

DE

DS

  • Last Week:
    • Meetings in Tucson
      • Pointing - M. Wagner, D. Thompson, etc.
      • Acquisition Camera Design & Implementation brainstorming- J. Kern, D. Thompson, John H., Christian
    • ARGOS pointing performance (first look)
    • Collimation model performance (Feb, March)
  • This week:
    • Pointing Log data evaluation - closer look to see if there might be problems with input data to TPoint
    • Collimation modelling equations (E-time prep)
    • ARGOS IE/CA analysis (from telemetry)
    • Begin Acquisition Camera design writeup for HW transition to Eng Team.

IS

Last week I worked mostly on the LBC script bug. Alsmot all was XSLT work

JA

jastier-dev now has 16 GB. Thanks Stephen!

Last week - Binocular planner, mostly copointing. Done except for one absorb case.

This week - make sure it can pass all tests from DS.

KS

Last week:
  • OVMS+TT meeting - the most interesting part for me was the discussion about synchronizing time between AO and OVMS
  • SVN diffs with OAC code to make sure current version is tagged before I check in the new HK code
  • Follow up of shutdown tasks - found humidity sensor calibration info and passed it on.
  • In Tucson on Friday:
    • configured 2 new spare iBootBars for LBC
    • meeting with Dan and Mike and then with John about what and where to put alarms in OAC upgrade
    • OVMS discussion with John - code update required and there's some fiber labeling on the mountain to check

This week:
  • OAC HK upgrade
    • Add client for setting warning thresholds
    • Start alarm handling - direct to EPICS
    • Clean up before checkin - error-handling, no test code, etc.
  • Meeting with Fernando for LBC controller upgrade
  • real-time code update for OVMS for engineering time (IT 7170)

MB

  • management OC, WPC, ISP interview
  • meeting Argos, OVMS+, LBTI post run
  • follow up VM SSD, NetBrain,SSP, AdSec PB plan
  • iteration and prototype git, rpm, Jenkins CI, rpm, yum server (ctd.)
  • issuetrak api - node.js scripts, API does not expose 'Last Activity Date', back to selenium
Next:
  • LBC controller upgrade [re]kickoff
  • responsibilities and assignments draft
  • techtalk#1 git//ci/rpm/kickstart

PK

  • Time To Live - Az/El axis working, problems with rotator code (will spend at least a part of this week on it)
  • DUT1 final testing, Saturday engineering
  • Missing PCS documentation??

SH

  • Packaging support scripts (e.g obsrunner.py, TCS simulator) -- what repo, and how to split them -- part of a common build?
  • RPM packaging web applications, with consideration of SE-Linux.
  • Using "dot" to draw complex diagrams.
  • turning on process accounting on obs computers with the eye on who is running what? Maybe go back over shell histories?

XZ

Last Week:
  • Meeting/diucssion about 4D related, OVMS+, powerBack plane and ARGOS
  • Sent detailed rip information to LBTI
  • Skip frame threshold discussion
  • Read the MG document and give feedback to MG
  • Discuss the flat apply issue on Adsec

This week:
  • Add log information for TT LUT
  • finish reading the MG document
  • write the low level communication code for PowerBack plane

YZ

  • General browse of observatory information.
  • Look at and study a bit of the SVN AO code, particularly trying to understand how the logging works.

Action items

  • NetBrain (David - schedule demo)
  • VM/SSD (David - followup meeting)
  • binoccular planning follow-up (Joe, Doug, Matthieu)
    • do we need the same resolution as TCS in OT? maybe the code complexity just Ok
  • Core A (Stephen - follow up, eng night?)
  • Argos (Andrew - s/w handover doc progress, schedule meeting)
  • Logging (Yang - requirements, design, concepts iteration - schedule meeting) - TCS lib (PK), other info in the Twiki (TBD)
  • SSP final draft (Matthieu)
  • techtalk#1 (Matthieu)
  • github user details, responsibilities, etc. - consult emails, meeting minutes, MB if needed (---)
  • file vacation and time off in time

Open discussion

  • pointing / collimation (DS) Thinking of putting up together a new camera/telescope to isolate mount vs other error source

-- %USERSIG{MatthieuBec}%

Comments

-- %BUBBLESIG{XianyuZhang - 2018-03-26}%
 
Topic revision: r22 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback